OpenStudy (anonymous):

what is the integral of 2/x^7+4/x^2

OpenStudy (blockcolder):

Rewrite that as 2x^(-7)+4x^(-2) and use the formula \[\int x^n \operatorname{d}\!x={x^{n+1} \over n+1}+C\]

OpenStudy (anonymous):

\[\int\limits_{?}^{?}x^{n}=x^{n+1}/n+1 +K n \neq -1\] n=-7

OpenStudy (campbell_st):

rewrite it in index form \[\int\limits2x^{-7} + 4x^{-2} dx = -2/6x^{-6} - 4x^{-1} + c\] or \[-1/(3x^6) - 4/x + c\]

OpenStudy (anonymous):

Thank you this gives me a better understanding